OpenStudy (anonymous):

To clear fractions from the equation , you could multiply each term on both sides by the number 8. True False

OpenStudy (anonymous):

hld on ima post the question

OpenStudy (anonymous):

\[\left(\begin{matrix}3 \\ n\end{matrix}\right)n-\left(\begin{matrix}11 \\ 2\end{matrix}\right)\]